# Low-Level and NUMA-Aware Optimization for High-Performance Quantum Simulation

### v1

@@ II-A Data Structure Enhancements
In the baseline QuEST implementation, complex amplitudes were stored in a structure-of-arrays (SoA) format. Each state-vector was represented by a ComplexArray object containing two separate arrays: one for all real parts and one for all imaginary parts. While this representation simplifies certain operations, it leads to inefficient memory access and poor cache utilization because accessing a complex amplitude requires touching two distant memory locations. We refactored this into an array-of-structures (AoS) layout, packing each complex number’s real and imaginary components contiguously in memory. This AoS format (illustrated in Figure 1) places the real and imaginary parts of each amplitude next to each other, improving spatial locality. As a result, when iterating through the state vector, each cache line fetched contains complete complex values, reducing cache misses and better exploiting hardware prefetching. We also ensure that the allocated array of complex numbers is aligned in memory, which facilitates aligned vector load/store operations in later optimizations. This transition from an AoS to a SoA allows for improved performance in both memory usage and data access patterns, particularly in high-performance contexts like quantum computing simulations, where large arrays of complex numbers are processed frequently.
@@ II-B NUMA-Aware Optimization
Modern multi-socket servers are NUMA: each CPU has faster access to memory on its local socket than on remote sockets. If left to the default OS policies, the memory for a process may be spread across NUMA nodes in a suboptimal way, and threads may migrate between sockets, incurring costly remote memory accesses. Often this is controlled in the HPC setting by pinning MPI processes to NUMA regions, but this approach is sub-optimal in QuEST, due to the additional memory buffers associated with MPI communication increasing the memory overhead. Thereby, in this section, we focus on two strategies to optimize memory placement and thread affinity on a dual-NUMA system.
@@ II-B1 Memory Allocation
By default, QuEST allocates the state-vector with the standard libc malloc, which on Linux follows a first-touch policy for NUMA (pages are allocated on the node where they are first accessed). This means that the OS is free to distribute pages across NUMA nodes arbitrarily, which may not be optimal on systems with multiple memory nodes, leading to performance penalties from remote memory accesses. To address this limitation, we introduce explicit control over memory placement via two NUMA-optimized strategies: (i) Local Allocation (V1), where the code first queries the free memory on each NUMA node (via numa_node_size64), then checks if the required memory can fit into a single node. If so, it performs an mmap allocation and sets a bind policy with mbind to restrict that allocation strictly to the chosen node. This ensures localized memory access – maximizing the benefit of a minimal memory latency. If the memory size exceeds what one node can provide, it gracefully falls back to partial allocation on the other node, while still enforcing page alignment. (ii) Split Allocation (V2), where we always split the allocated memory evenly across two nodes from the start. After computing halfSize (and ensuring that it remains at least one page in size), we allocate the entire region via mmap, then explicitly bind the first half of the region to Node 0 and the second half to Node 1. This approach ensures a balanced distribution of memory which yields higher total memory bandwidth. As in V1, the routine first rounds the requested buffer size up to an exact multiple of the system page size, allocates it with mmap, and verifies that the returned pointer is 64-byte aligned – guaranteeing efficient AVX-512 loads and stores. Both strategies confirm that the base address still starts on a page boundary, eliminating sub-page fragmentation that could otherwise hamper memory-access performance. Figure 2 compares the available allocation strategies schematically. These methods integrate seamlessly within the QuEST environment, allowing the user to switch between Default, NUMA V1, or NUMA V2 at compile or runtime, depending on the specific memory requirements and system architecture.
@@ II-B2 Thread Affinity
Locality‐aware task scheduling is essential in NUMA architectures because it ensures that threads run on the memory node where their data reside, thereby minimizing costly remote accesses. In our approach, we dynamically calculate task chunks and bind threads to specific NUMA nodes. Instead of relying on the OS scheduler, we bind half of the threads to the first socket and the other half to the second socket – aligned with layout used in allocation V2 – see Algorithm 2. Moreover, we divide the state-vector workload such that each thread mostly works on the portion of the state-vector residing in its local node’s memory. In our two-node system, threads process the lower half of the amplitude array (allocated on node 0) and threads process the upper half (allocated on node 1). This manual partitioning and pinning maximizes data locality for most workloads and prevents idle resources when tasks are not perfectly divisible. Together with optimizations such as loop unrolling and cache‐friendly data layouts, this strategy significantly lowers memory‐access latency and boosts overall performance. Figure 3 evaluates the performance impact of NUMA-aware optimizations on a sequence of 50 Hadamard gates, profiled with Intel PCM NUMA [19]. The Y-axes show the DRAM accesses as seen by different threads (on the X-axes), for different qubits. Configuration (a) employs the baseline QuEST implementation, leading to suboptimal/scattered accesses across nodes. Configuration (b) introduces NUMA-aware memory allocation while still using the baseline Hadamard gate, reducing remote memory accesses and improving memory locality by preferentially utilizing local memory. Configuration (c) further refines this approach utilising a locality‐sensitive scheduler. As a result, when the working set fits entirely within a single NUMA node – up to 34 qubits on our test system – all DRAM accesses remain local, effectively minimizing remote memory traffic. When total memory demand spans multiple nodes (35 qubits), the state-vector spans NUMA nodes, resulting in a mix of local and remote DRAM accesses, thereby still minimizing cross‐node traffic. Moreover, similar improvements can be achieved by applying environment settings (e.g., OMP_PROC_BIND, OMP_PLACES, GOMP_CPU_AFFINITY) that enforce core affinity, confirming the critical role of thread binding in reducing latency.
@@ II-C Instruction-level Tuning
With the data layout and NUMA strategy in place, we further optimize the innermost loop of the simulation using advanced instruction-level techniques.
@@ II-C1 Vectorization with AVX-512 Intrinsics
Vectorization allows a processor to more efficiently exploit data-parallelism, by performing identical instructions on different data in a single issue cycle. This often leads to performance gains when applications exhibit this pattern of having the same operation to many operands. In the case of state-vector simulation, the core operation is a matrix-vector multiplication, which exhibits enormous data parallelism, given the known regularity of the matrix sparsity pattern. However, enabling compiler optimizations at the compiler is insufficient to see any effect. This is due to how the code is written. In particular, the core of the gate loops in QuEST v3.7 rely heavily on bitwise operations, in particular masking and shifting, to identify which iterations to skip in controlled gates, and to calculate accessed indices from the iteration index. This pattern that does not translate well to vectorised instructions. Additionally, even when vectorising is working as expected and instructions are issued out faster and more efficiently, the memory bandwidth bottleneck still prevents translating these gains in faster gate execution times. In our case, we explicitly write our optimized kernels using intrinsic AVX-512 instructions, which permit simultaneous operations on eight double-precision floating-point values (4 complex amplitudes). Using aligned vector loads and stores (via _mm512_load_pd and _mm512_store_pd), the code minimizes memory transfer overhead. Critical arithmetic is further optimized using FMA instructions (_mm512_fmadd_pd and _mm512_fmsub_pd), which fuse multiplication with addition or subtraction operations common in quantum gate updates, thus further reducing the instruction count and improving numerical accuracy.
@@ II-C2 Loop Unrolling and Pointer Restriction
Aggressive loop unrolling is employed – processing blocks of 16 iterations at a time – to diminish the overhead of loop control and enhance instruction-level parallelism to the CPU. The use of the restrict qualifier further assists the compiler by confirming non-aliasing pointers, allowing more aggressive optimizations during vectorization.
@@ II-C4 Memory Prefetching
To further reduce memory access latency, for each group of 4 complex numbers processed, the code explicitly employs prefetching via _mm_prefetch to load future cache lines in advance, reducing cache miss penalties when processing large blocks of data. Our kernel processes 8 doubles at a time, hence prefetching data 16 and 32 elements ahead ensures that the subsequent iterations’ data is already loaded into the L1 cache.
@@ II-C5 Scalar Fallback for Complete Coverage
These optimizations are illustrated in pseudocode: Algorithm 1 shows the baseline Hadamard transformation incorporating the new AoS, while Algorithm 2 shows our fully optimised kernel. This implementation is engineered to maximize throughput on modern multi-core architectures through a series of low-level optimizations presented in Section II.
@@ III Benchmark Results and Analysis
The experiments were conducted on a dual-socket Intel® Xeon® Gold 6230R system (x86_64) featuring 52 physical cores (26 per socket), with hyperthreading disabled to ensure consistent core performance. Each core includes 32KiB instruction and 32KiB data L1 cache, and 1 MiB L2 cache. Each socket includes 35.75MiB L3 cache, for a total of 71.5 MiB amongst the two sockets. The system is configured with two NUMA nodes, each offering approximately 385 GB of local memory. Notably, local memory latency is about 100 ns, while remote memory latency reaches 150 ns, as measured using Intel’s MLC, highlighting the importance of NUMA-aware optimizations. To ensure stable and reproducible performance results, we disabled Intel Turbo Boost and turned off CPU frequency scaling, preventing transient clock speed variations from influencing the benchmarks. All experiments were performed on a Linux environment, and the code was compiled in Release mode with GCC 9.3.0 (C99, C++98) via CMake 3.7, employing aggressive optimization flags (e.g., -O3) to fully leverage the hardware’s performance capabilities. Moreover, for consistent benchmarking, we utilized Google Benchmark integrated within the Yao framework [20], where
to enhance statistical reliability each benchmark operation (i.e., an individual gate or a composite circuit) was repeated ten times and the minimum simulation time is reported.
@@ III-A Single- and Two-Qubit Operations
Figure 6(a) reports the absolute execution time per gate for a representative single-qubit gate (Hadamard). The baseline QuEST results are compared against multiple optimization strategies, and the effectiveness of these optimizations is more evident in Figure 6(b), which shows the corresponding speedups relative to the baseline. The overall performance gains observed across our optimizations indicate the clear benefits of progressively more aggressive low-level tuning. Reordering the data into an AoS alone yields a modest speedup of 1.2× over the baseline. Adding AVX2 vectorization dramatically increases the speedup – 5× faster than baseline. When we switch from AVX2 to AVX-512 (with the AoS layout), the speedup improves further to 5.5×. The best performance is achieved by the most aggressive optimization (AoS + AVX-512 + FMA + loop unrolling + prefetching), see Section II-C and Algorithm 2, reaching a speedup in the vicinity of 6× compared to the baseline. These gains reflect how each low-level improvement builds upon the previous ones and can be directly attributed to the underlying memory-access patterns of quantum operations. The restructured AoS format enhances spatial locality, which in turn enables efficient SIMD loading of data.
@@
Building on the optimizations introduced in Figure 6(b), Figure 7 extends this analysis to compare the three gate types (Hadamard, Pauli-, and CNOT) using the fully optimized kernel under the two NUMA memory allocation strategies (see Section II-B1). V1 (node‐0-first) remains faster than V2 (split allocation) up to the penultimate qubit size. In this region, the memory footprint fits comfortably within a single NUMA node, so the overheads of remote accesses are minimal, and local allocation yields lower latency. However, once we reach the largest qubit size (where one node’s memory no longer suffices), V1 is forced to allocate data on the remote node as well, causing its performance to degrade. At this point, V2 (which evenly distributes memory across both nodes) achieves comparable or better results by balancing the load between nodes. Another key observation is that speed-up improvements differ notably among the three gate types. For Hadamard and Pauli- operations, we achieve higher speedups (up to 5.5-6.5×) over the baseline compared to CNOT (which peaks around 4.5×). This behavior is closely tied to the memory‐access patterns depicted in Figures 5. In contrast to the single-qubit gates, which involve uniform, independent rotations on all amplitudes (ideal for vectorized execution), CNOT involves controlled operations, where certain branches of the matrix‐vector multiply are skipped depending on the control qubit’s state – only the amplitudes corresponding to a control bit of 1 are updated. This conditional behavior introduces branches in the matrix‐vector multiply, as certain computational paths are effectively skipped. This branching logic and the more irregular, conditional nature of the data manipulation reduces opportunities for continuous SIMD operations and hampers consistent cache reuse. Consequently, although CNOT still benefits from multi‐threading and low-level tuning strategies, it does not achieve speedups as high as the simpler single‐qubit operations.

### v2

@@ II-A Data Structure Enhancements
In the baseline QuEST v3.7.0 implementation, complex amplitudes are stored in a structure-of-arrays (SoA) format: a ComplexArray with two separate arrays, one for all real parts and one for all imaginary parts. While this representation simplifies certain operations, it leads to inefficient memory access and poor cache utilization because accessing a single complex amplitude requires touching two distant memory locations. We refactored this into an array-of-structures (AoS) layout, packing each amplitude’s real and imaginary components contiguously in memory. This AoS format (illustrated in Figure 1) places the real and imaginary parts of each amplitude next to each other, improving spatial locality. As a result, when iterating through the state vector, each cache line fetched contains complete complex values, reducing cache misses and better exploiting hardware prefetching. We also ensure that the allocated array of complex numbers is aligned in memory (64-byte alignment), which facilitates aligned vector load/store operations in later optimizations. This transition from an AoS to a SoA improves effective cache use and SIMD efficiency in large-scale quantum circuit simulations, where large arrays of complex numbers are processed frequently.
@@ II-B NUMA-Aware Optimization
Modern multi-socket servers employ NUMA, where each socket has local memory with higher bandwidth and lower latency than remote memory connected through inter-socket links such as UPI or Infinity Fabric. Without NUMA-awareness–i.e., if left to the default OS policies–the memory of a process may be spread across sockets in a suboptimal way, and threads may migrate between nodes, leading to costly remote memory accesses. In multi-threaded simulations such as QuEST, these effects can severely limit scaling efficiency if threads frequently access data residing on another socket. While in distributed-memory (MPI-based) settings this issue is often mitigated by binding MPI ranks to NUMA regions, our focus here is on single-process, multi-threaded execution where such process-level pinning is not directly applicable. Therefore, in this section, we employ two complementary strategies to optimize memory placement and thread affinity on a dual- (multi-) NUMA system.
@@ II-B1 Memory Allocation
By default, QuEST allocates the state vector with the standard libc malloc, which on Linux follows a first-touch policy for NUMA (pages are allocated on the node where they are first accessed). This means that the OS is free to distribute pages across NUMA nodes arbitrarily, which may not be optimal on systems with multiple memory nodes, leading to performance penalties from remote memory accesses. To address this limitation, we introduce explicit control over memory placement via two NUMA-optimized strategies: (i) Local Allocation (V1), where the code first queries the free memory on each NUMA node (via numa_node_size64), then checks if the required memory can fit into a single node. If so, it performs an mmap allocation and sets a bind policy with mbind to restrict that allocation strictly to the chosen node. This ensures localized memory access, maximizing the benefit of low-latency local memory. If the memory size exceeds what one node can provide, it gracefully falls back to partial allocation on the other node, while still enforcing page alignment. (ii) Split Allocation (V2), where we always split the allocated memory evenly across two nodes from the start. After computing halfSize (and ensuring that it remains at least one page in size), we allocate the entire region via mmap, then explicitly bind the first half of the region to Node 0 and the second half to Node 1. This approach ensures a balanced distribution of memory which yields higher total memory bandwidth. As in V1, the routine first rounds the requested buffer size up to an exact multiple of the system page size, allocates it with mmap, and verifies that the returned pointer is 64-byte aligned–guaranteeing efficient AVX-512 loads and stores. Both strategies confirm that the base address still starts on a page boundary, eliminating sub-page fragmentation that could otherwise hamper memory-access performance. Figure 2 compares the available allocation strategies schematically. These methods integrate seamlessly within the QuEST environment, allowing the user to switch between Default, NUMA V1, or NUMA V2 at compile or runtime, depending on the specific memory requirements and system architecture.
@@ II-B2 Thread Affinity
Locality‐aware task scheduling is essential in NUMA architectures, as it ensures that threads execute on the memory node where their data reside, thereby minimizing costly remote accesses. In our approach, we dynamically calculate task chunks and bind threads explicitly to NUMA nodes rather than relying on the OS’s scheduler. Half of the threads are pinned to the first socket, and the other half to the second socket–aligned with the layout used in allocation V2 (see Algorithm 9). The state vector workload is divided accordingly so that each thread operates primarily on the portion of the state vector stored in its local node’s memory. In our two-socket system, threads process the lower half of the amplitude array (allocated on node 0), while threads process the upper half (allocated on node 1). Algorithm 2 illustrates the NUMA-aware thread binding strategy used in our implementation. During initialization, each OpenMP thread determines its global ID and binds itself to the corresponding NUMA node, ensuring memory locality by keeping both the working set and execution context within the same NUMA domain and thereby reducing cross-node memory latency.
This manual partitioning and pinning strategy maximizes data locality for most workloads and prevents idle resources when tasks are not perfectly divisible. Combined with optimizations such as loop unrolling and cache‐friendly data layouts, it significantly lowers memory‐access latency and boosts overall performance.
Figure 5 evaluates the impact of NUMA-aware optimizations on a sequence of 50 Hadamard gates, profiled with Intel PCM NUMA [21]. The Y-axis shows DRAM accesses observed by individual threads (on the X-axis) for different target qubits. Configuration (a) represents the baseline QuEST implementation, exhibiting scattered and suboptimal memory accesses across nodes. Configuration (b) introduces NUMA-aware memory allocation while retaining the baseline Hadamard kernel, thereby improving locality by favoring accesses to local memory. Configuration (c) further refines this approach by employing a locality‐sensitive scheduler. As a result, when the working set fits entirely within a single NUMA node–up to 34 qubits on our test system–all DRAM accesses remain local, effectively minimizing remote memory traffic. When the state vector spans multiple NUMA nodes (e.g., at 35 qubits), a mix of local and remote accesses occurs, yet cross-node traffic remains significantly reduced. Comparable improvements can also be achieved using environment variables (e.g., OMP_PROC_BIND, OMP_PLACES, GOMP_CPU_AFFINITY) to enforce core affinity, confirming the critical role of thread binding in reducing latency. However, manually toggling and tuning these environment variables can be error-prone and tedious, especially across systems with different NUMA topologies. In contrast, our implementation performs this binding automatically based on the detected NUMA topology, while still allowing users to override or adjust the mapping policy at compile time or runtime if needed.
@@ II-C Instruction-level Tuning
With the optimized data layout and NUMA strategy established, we further enhance the innermost simulation loops through instruction-level optimizations that exploit fine-grained hardware parallelism.
@@ II-C1 Vectorization with AVX-512 Intrinsics
Vectorization enables a processor to exploit data-level parallelism by performing identical operations on multiple data elements within a single instruction cycle. This often yields substantial performance gains for applications that repeatedly apply the same operation across large datasets. In state vector simulation, the core operation–a matrix-vector multiplication–exhibits strong data parallelism due to the regular and sparse structure of quantum gate matrices.
However, relying solely on compiler auto-vectorization is insufficient to achieve meaningful speedups because of the way the original QuEST code is written. In particular, the gate-application loops in QuEST v3.7.0 rely heavily on bitwise operations, in particular masking and shifting, to identify which iterations to skip for controlled gates and to compute the corresponding memory indices from the iteration index. Such irregular access patterns do not translate efficiently into vectorized instructions. Moreover, even when vectorization succeeds and instructions are issued at higher throughput, overall performance remains limited by memory bandwidth, which prevents these gains from directly translating into faster gate execution times.
In our implementation, we explicitly write the optimized kernels using intrinsic AVX-512 instructions, which perform simultaneous operations on eight double-precision floating-point values (corresponding to four complex amplitudes). By using aligned vector loads and stores (via _mm512_load_pd and _mm512_store_pd), the code minimizes memory-transfer overhead. Arithmetic operations are further optimized with FMA instructions, which combine multiplication and addition (or subtraction) in a single operation–common in quantum gate updates–thereby further reducing the instruction count and improving numerical accuracy.
Algorithm 3 outlines the high-level AVX-512 vectorization strategy used for the Hadamard kernel. Each loop iteration processes four amplitude pairs simultaneously (eight complex double-precision numbers). The intrinsic operations _mm512_load_pd, _mm512_fmadd_pd, and _mm512_fmsub_pd are abstracted here as high-level vector load, fused multiply-add, and store operations, enabling efficient SIMD parallelism with minimal memory latency.
While Algorithm 3 presents the vectorization at a high level, the actual compute loop is implemented as a dedicated AVX-512 mini-kernel that fuses multiplication and addition/subtraction within a single instruction for maximum throughput. This mini-kernel is repeatedly invoked from the main optimized kernel (Algorithm 9).
@@ II-C2 Loop Unrolling and Pointer Restriction
Aggressive loop unrolling is employed–processing blocks of 16 iterations per step–to reduce the overhead of loop control and enhance instruction-level parallelism on the CPU. The use of the restrict qualifier further assists the compiler by confirming non-aliasing pointers, enabling more aggressive vectorization and other low-level optimizations, as illustrated in Algorithm 5.
@@ II-C3 Memory Optimization and NUMA Awareness
In addition to thread binding, we explicitly partition the state vector between NUMA nodes to ensure memory is allocated close to the threads that will access it. As illustrated in Algorithm 6, the state vector is divided into contiguous chunks, with each chunk allocated on the memory node corresponding to a specific NUMA domain. Threads pinned to a node process only their local chunk of the state vector, ensuring that nearly all memory accesses remain local. This NUMA-aware partitioning works synergistically with thread affinity (Algorithm 2): the former places data where it belongs, while the latter ensures computation happens near the data. Together, these reduce remote DRAM accesses and improve effective bandwidth utilization. When the state vector fits within a single NUMA domain, all accesses are local; as the problem scales across nodes, this partitioning strategy minimizes—but cannot entirely eliminate—remote memory traffic.
@@ II-C4 Memory Prefetching
To further reduce memory access latency, for each group of 4 complex numbers processed, the code explicitly employs prefetching via _mm_prefetch to load future cache lines in advance, reducing cache miss penalties when processing large blocks of data. This is critical for wide vector loads (_mm512_load_pd) to sustain bandwidth. Our kernel processes eight doubles at a time, hence prefetching data 16 and 32 elements ahead ensures that the subsequent iterations’ data is already loaded into the L1 cache.
@@ II-C5 Scalar Fallback for Complete Coverage
These optimizations are illustrated in pseudocode: Algorithm 1 shows the baseline Hadamard transformation incorporating the new AoS, while Algorithm 9 shows our fully optimised kernel. This implementation is engineered to maximize throughput on modern multi-core architectures through a series of low-level optimizations222Our AVX-512 implementation goes beyond simply widening AVX2 with FMA instructions. We employ a 512-bit aligned SoA layout, unroll loops in blocks of 16 iterations (corresponding to 4 complex amplitudes), and apply fixed-distance prefetching (e.g., +16, +32 elements) to reduce L1 cache latency. These elements, combined with carefully hand-written FMA-based vector kernels, significantly improve cache utilization and instruction throughput compared to the AVX2-AoS baseline. presented in Section II.
@@ III Benchmark Results and Analysis
The experiments were conducted on a dual-socket Intel® Xeon® Gold 6230R system (x86_64) featuring 52 physical cores (26 per socket), with hyperthreading disabled to ensure consistent core performance. Each core includes 32 KiB of instruction and 32 KiB of data L1 cache, and 1 MiB of L2 cache. Each socket includes 35.75MiB L3 cache, for a total of 71.5 MiB amongst the two sockets. The system is configured with two NUMA nodes, each offering approximately 385 GB of local memory. Notably, local memory latency is about 100 ns, while remote memory latency reaches 150 ns, as measured using Intel’s MLC, highlighting the importance of NUMA-aware optimizations. To ensure stable and reproducible performance results, we disabled Intel Turbo Boost and turned off CPU frequency scaling, preventing transient clock speed variations from influencing the benchmarks. All experiments were performed on a Linux environment, and the code was compiled in Release mode with GCC 9.3.0 (C99, C++98) via CMake 3.7, employing aggressive optimization flags (e.g., -O3) to fully leverage the hardware’s performance capabilities. Moreover, for consistent benchmarking, we utilized Google Benchmark integrated within the Yao framework [22], where to enhance statistical reliability each benchmark operation (i.e., an individual gate or a composite circuit) was repeated ten times and the minimum simulation time is reported.
@@ III-A Single- and Two-Qubit Operations
Figure 6(a) reports the absolute execution time per gate for a representative single-qubit gate (Hadamard). The baseline (default v3.7.0 build) QuEST results are compared against multiple optimization strategies, and the effectiveness of these optimizations is more evident in Figure 6(b), which shows the corresponding speedups relative to the baseline. The overall performance gains observed across our optimizations indicate the clear benefits of progressively more aggressive low-level tuning. Reordering the data into an AoS alone yields a modest speedup of 1.2 over the baseline. Adding AVX2 vectorization dramatically increases the speedup–5 faster than baseline. When we switch from AVX2 to AVX-512 (with the AoS layout), the speedup improves further to 5.5. The best performance is achieved by the most aggressive optimization (AoS + AVX-512 + FMA + loop unrolling + prefetching), see Section II-C, and Algorithm 9, reaching a speedup in the vicinity of 6 compared to the baseline. These gains reflect how each low-level improvement builds upon the previous ones and can be directly attributed to the underlying memory-access patterns of quantum operations. The restructured AoS format enhances spatial locality, which in turn enables efficient SIMD loading of data.
@@
Building on the optimizations introduced in Figure 6(b), Figure 7 extends this analysis to compare the three gate types (Hadamard, Pauli-, and CNOT) using the fully optimized kernel under the two NUMA memory allocation strategies (see Section II-B1). V1 (node‐0-first) remains faster than V2 (split allocation) up to the penultimate qubit size. In this region, the memory footprint fits comfortably within a single NUMA node, so the overheads of remote accesses are minimal, and local allocation yields lower latency. However, once we reach the largest qubit size (where one node’s memory no longer suffices), V1 is forced to allocate data on the remote node as well, causing its performance to degrade. At this point, V2 (which evenly distributes memory across both nodes) achieves comparable or better results by balancing the load between nodes. Another key observation is that speed-up improvements differ notably among the three gate types. For Hadamard and Pauli- operations, we achieve higher speedups (up to 5.5-6.5) over the baseline compared to CNOT (which peaks around 4.5). This behavior is closely tied to the memory‐access patterns depicted in Figures 4. In contrast to the single-qubit gates, which involve uniform, independent rotations on all amplitudes (ideal for vectorized execution), CNOT involves controlled operations, where certain branches of the matrix‐vector multiply are skipped depending on the control qubit’s state–only the amplitudes corresponding to a control bit of 1 are updated. This conditional behavior introduces branches in the matrix‐vector multiply, as certain computational paths are effectively skipped. This branching logic and the more irregular, conditional nature of the data manipulation reduces opportunities for continuous SIMD operations and hampers consistent cache reuse. Consequently, although CNOT still benefits from multi‐threading and low-level tuning strategies, it does not achieve speedups as high as the simpler single‐qubit operations.